A cork of density 0.5gcm^-3 floats on a calm swimming pool The fraction of the Cork`s volume which is under water

Grade:

1 Answers

suvidhi mehta
6 years ago
let x be length of cork immersed
and l be its total length
and A be its cross sectional area
for floating condition , weight = bouyant force
l*A*(density of cork)g = (density of liquid)*x*A*g
fraction=x/l=density of cork/density of liquid=0.5/1 = 0.5
hope its correct and helps u